# HyperMoE: Towards Better Mixture of Experts via Transferring Among Experts

### v1

@@ 5.1 Mixture of Expert
Shazeer et al., (2017) introduces Mixture-of-Expert layers for LSTM language modeling and machine translation. These layers are inserted between the standard layers of the LSTM model. Subsequent work primarily builds on Transformers, where expert layers often replace dense layers. Lepikhin et al., (2021) first introduces MoE layers into Transformers and studies them in the context of machine translation. With the release of Gshard (Lepikhin et al.,, 2021) and Switch Transformer (Fedus et al.,, 2022), MoE models are scaled up to new heights by introducing thousands of small-scale experts. In terms of routing, Shazeer et al., (2017) use routing to the top k experts out of k > 1. Hazimeh et al., (2021) propose DSelect-k, a smoothed version of the top-k routing algorithm that improves upon standard top-k routing. Fedus et al., (2022), Clark et al., (2022) and Xue et al., (2023) demonstrate that top-1 routing can also achieve competitive results. Hash Layer (Roller et al., 2021b, ) and StableMoE (Dai et al.,, 2022) employ fixed routing strategies for more stable routing and training. Zhou et al., (2022) propose an expert selection routing strategy where each token can be assigned to a different number of experts. Rajbhandari et al., (2022) and Dai et al., (2024) isolate general knowledge from experts using shared experts from engineering and algorithm perspectives, respectively, to promote expert specialization.
@@ 5.2 HyperNetwork
Hypernetworks (Ha et al.,, 2017) are widely used in multi-task learning due to their ability to avoid negative interference of corresponding modules by soft parameter sharing and generating module parameters conditioned on the shared parameters. The most common approach usually takes task (Karimi Mahabadi et al.,, 2021; Zhao et al.,, 2023) or language embeddings (Üstün et al.,, 2020; Baziotis et al.,, 2022) as contextual information to generate corresponding module parameters, such as adapter layers (Üstün et al.,, 2020; Ansell et al.,, 2021; Karimi Mahabadi et al.,, 2021), classifier heads (Ponti et al.,, 2021), and continuous prompts (He et al.,, 2022). In addition, hypernetwork-based approaches have also been very successful in zero-shot and few-shot scenarios (Deb et al.,, 2022; Phang et al.,, 2023; Ivison et al.,, 2023; Jin et al.,, 2024).
In the field of NLP, hypernetworks are mainly used to improve the generalization (Volk et al.,, 2022; Zhang et al.,, 2023) and applicability (Wullach et al.,, 2022; He et al.,, 2022; Tan et al.,, 2023) of dense models. Our work explores the integration of hypernetworks with sparse MoE. We propose to input the expert selection status of tokens as information into the hypernetwork and generate module parameters that correspond to the respective tokens. To the best of our knowledge, this is the first time that hypernetworks have been introduced in the MoE structure, which extends the application scope of hypernetworks and provides new insights for knowledge transferring in MoE.
@@ 6 Conclusion
Despite our work has demonstrated strong experimental results, there are several limitations: (1) In this work, we utilize end-to-end training to learn expert embeddings. Incorporating prior knowledge, such as expert weights, into the embedding learning process may improve efficiency and performance. We will improve upon this in future work. (2) We insert HyperExpert into the expert layer of MoE in parallel. This incurs additional computational overhead. Mitigating this issue could be achieved by employing some parameter-efficient methods (such as LoRA (Hu et al.,, 2022) and (IA)${}^{3}$ (Liu et al.,, 2022)) to insert HyperExpert into MoE. (3) Current experiments mainly focus on fine-tuning the pre-trained MoE model. Utilizing our proposed method to train a large-scale MoE from scratch will be the emphasis of our future work.

### v3

@@ 5.1 Mixture of Expert
Shazeer et al. (2017) introduces Mixture-of-Expert layers for LSTM language modeling and machine translation. These layers are inserted between the standard layers of the LSTM model. Subsequent work primarily builds on Transformers, where expert layers often replace dense layers. Lepikhin et al. (2021) first introduces MoE layers into Transformers and studies them in the context of machine translation. With the release of Gshard (Lepikhin et al., 2021) and Switch Transformer (Fedus et al., 2022), MoE models are scaled up to new heights by introducing thousands of small-scale experts. In terms of routing, Shazeer et al. (2017) use routing to the top k experts out of k > 1. Hazimeh et al. (2021) propose DSelect-k, a smoothed version of the top-k routing algorithm that improves upon standard top-k routing. Fedus et al. (2022), Clark et al. (2022) and Xue et al. (2023) demonstrate that top-1 routing can also achieve competitive results. Hash Layer (Roller et al., 2021b) and StableMoE (Dai et al., 2022) employ fixed routing strategies for more stable routing and training. Zhou et al. (2022) propose an expert selection routing strategy where each token can be assigned to a different number of experts. Qiu et al. (2023) demonstrates that a standard LM can be fine-tuned as its Mix of Experts (MoE) counterpart, effectively improving downstream in-domain and out-of-domain generalization capabilities. Rajbhandari et al. (2022) and Dai et al. (2024) isolate general knowledge from experts using shared experts from engineering and algorithm perspectives, respectively, to promote expert specialization.
@@ 5.2 HyperNetwork
Hypernetworks (Ha et al., 2017) are widely used in multi-task learning due to their ability to avoid negative interference of corresponding modules by soft parameter sharing and generating module parameters conditioned on the shared parameters. The most common approach usually takes task (Karimi Mahabadi et al., 2021; Zhao et al., 2023) or language embeddings (Üstün et al., 2020; Baziotis et al., 2022) as contextual information to generate corresponding module parameters, such as adapter layers (Üstün et al., 2020; Ansell et al., 2021; Karimi Mahabadi et al., 2021), classifier heads (Ponti et al., 2021), and continuous prompts (He et al., 2022). In addition, hypernetwork-based approaches have also been very successful in zero-shot and few-shot scenarios (Deb et al., 2022; Phang et al., 2023; Ivison et al., 2023).
In the field of NLP, hypernetworks are mainly used to improve the generalization (Volk et al., 2022; Zhang et al., 2023) and applicability (Wullach et al., 2022; He et al., 2022; Tan et al., 2023) of dense models. HyperRouter (Do et al., 2023) chooses to solve the routing allocation problem in SMoE methods by using a fixed HyperNetwork to generate routing weights. Our work explores the integration of hypernetworks with sparse MoE. We propose to input the expert selection status of tokens as information into the hypernetwork and generate module parameters that correspond to the respective tokens. To the best of our knowledge, this is the first time that hypernetworks have been introduced in the MoE structure, this is the first time HyperNetwork has been applied to the expert structure in MoEs, introducing inductive bias through the hypernetwork: information from unselected experts is reused, which extends the application scope of hypernetworks and provides new insights for knowledge transferring in MoE.
@@ 6 Conclusion
Despite our work has demonstrated strong experimental results, there are several limitations: (1) In this work, we utilize end-to-end training to learn expert embeddings. Incorporating prior knowledge, such as expert weights, into the embedding learning process may improve efficiency and performance. We will improve upon this in future work. (2) We insert HyperExpert into the expert layer of MoE in parallel. This incurs additional computational overhead. Mitigating this issue could be achieved by employing some parameter-efficient methods (such as LoRA (Hu et al., 2022) and (IA)3 (Liu et al., 2022)) to insert HyperExpert into MoE. (3) Current experiments mainly focus on fine-tuning the pre-trained MoE model. Utilizing our proposed method to train a large-scale MoE from scratch will be the emphasis of our future work.
@@ Appendix A Depthwise Separable Convolutional Networks Details
Appendix B Additional Results
Here, We conducted additional experiments to demonstrate the superiority of our method over similar methods. The experiment is based on SwitchTransformer and GPT2-MoE, with all methods having the same number of parameters and experts. Moreover, we trained GPT2-MoE from scratch on WikiText-103, and the experimental results are as shown in Table 7.
Method SQuAD CNN WikiText-2 WikiText-103 (from scratch)
MoE 83.01 19.75 21.71 22.0
HyperRouter 78.77 15.56 21.69 22.21
HyperMoE 83.51 20.21 21.49 21.81
Table 7: Results of the comparison with HyperRouter on 4 datasets(SQuAD, CNN, WikiText-2, WikiText-103). We conduct pre-training experiments on WikiText-103, while the other three datasets are used for fine-tuning experiments.
The results show that our method achieves the best performance both scratch training and finetuning. The way HyperRouter generates router weights causes the model to lose its original pre-trained weights, making HyperRouter ineffective in fine-tuning scenarios. The reason HyperRouter’s performance is inconsistent with the original results under the scratch training scenario might be because we fixed the number of experts chosen during the training process, instead of gradually increasing the number of chosen experts during training as in the original setup. In the original setting, the computational cost of HyperRouter is significantly higher than the method with a fixed number of experts. For a fair comparison, we chose the same expert selection setting. Experimental results show that HyperMoE can efficiently transfer knowledge between different experts in MoE when choosing the same number of experts.